OpenStudy (anonymous):

What is the third perfect number? Please help me understand how it works. thnx

hero (hero):

1^2 = 1 2^2 = 4 3^2 = 9 The third perfect number is 9? Did you mean perfect SQUARES?

OpenStudy (anonymous):

ohh, no, ummm, like the first number is 6, the second is 28, and idk what the third is...plus idk how to get the number

OpenStudy (anonymous):

Aperfect number is a positive integer that is equal to the sum of its proper positive divisors. Mathematically, \( \sigma_1(N)=2N \) \(\sigma_1() \) is the divisor function ( http://en.wikipedia.org/wiki/Divisor_function) Here is the list of perfect numbers in OEIS: http://oeis.org/A000396

hero (hero):

Never heard of it

hero (hero):

And fool clearly wiki'd it

OpenStudy (anonymous):

A year back I wrote a very optimized perfect number generator.

OpenStudy (anonymous):

whoa, did you code it all?

OpenStudy (anonymous):

Unfortunately it's a part of an algorithm contest, so I can't share the code here. But I can give some hints if you want :)

OpenStudy (anonymous):

Hero, I don't really need to wike perfect numbers :)

OpenStudy (anonymous):

Yes please, any that are available :) I just started coding a few days ago and I'm working through the Python textbook. Just started my first more complex code (finding slope and y-intercept)still very simple, but I'd appreciate it

OpenStudy (anonymous):

Okay first there are no odd perfect numbers (this is not proven yet) so you can clearly ignore the odd numbers.

OpenStudy (anonymous):

ok

hero (hero):

Why am I just now hearing about perfect numbers? In what course do you usually learn about them?

hero (hero):

14yamaka, what course are you currently taking?

OpenStudy (anonymous):

Now, you need to have a very fast \( \sigma_1(N) \) for that you need to have a fast optimized prime generator. This is actually deep I have used Miller Rabin algorithm.

OpenStudy (anonymous):

@Hero: Elementary Number theory.

OpenStudy (anonymous):

14yamaka, I would probably advice you to be satisfied with typical brute-force at this stage :)

OpenStudy (anonymous):

*This is actually deep, I have used Miller Rabin algorithm + sieve.

OpenStudy (anonymous):

wow, ok, I gotta read up on this Miller Rabin algorithm. and sieve. Thanks for the explanation. Both of you :)

OpenStudy (anonymous):

Fain I help :)

hero (hero):

I didn't provide anything, lol FFM provided everything